ADNAN MUHAMMED

Grade 12,

two circles intersect each other at the points P and Q. If AB and AC are tangents to the two circles from a point A on the line joining P and Q, show that AB=AC

two circles intersect each other at the points P and Q.
If AB and AC are tangents to the two circles from a point A on the line joining P and Q, show that AB=AC

Grade:11

1 Answers

Madhur Nanhorya
6 years ago
Now AB^2=AP×PQ______(1)Also AC^2=AP×PQ_______(2)From eqn 1 and 2 we get AB^2=AC^2Therefore AB=ACDHence Proved.